I am struggling with this word problem:
Business and finance. The revenue for a sandwich shop is directly proportional to its advertising budget. When the owner spent $2000 a month on advertising, the revenue was $120,000. If the revenue is now $180,000, how much is the owner spending on advertising?

Answer by sarah_adam(201)

First, we set up our general equation. Because y is directly proportional to x, we have: y = cx
where c is the constant of proportionality. In other words, when x goes up, y goes up, and when x goes down, y goes down.
so here lets assume when x = 2000 , y = 120,000
so we need to find the value of x when y = 180,000
y = cx
120,000 = c (2000)
c = 120,000/2000 = 60
therefore when c = 60 and y = 180,000
180,000 = 60 (x)
x = 180,000/60 = 3,000
When the revenue is 180,000 the amount spent on advertising is $3000